the ratio of interest between the compound and simple interest for two years on a sum of money to that after three years on the same sum, at the same rate of interest, is 11: 37. What will be the rate of interest?
1
Expert's answer
2020-05-21T17:36:47-0400

Simple interest after two years is "P*2R" ,

after three years "P*3R" , where P is the sum of money, R - interest rate.


Compound interest after two years is "P*(1+R)^2-P"

after three years "P*(1+R)^3-P"


Difference between compound and simple interest on certain sum of money after two years is:

"P*(1+R)^2-P-P*2R=P*R^2"


Difference between compound and simple interest on certain sum of money after three years is:

"P*(1+R)^3-P-P*3R=3P*R^2+P*R^3"

Therefore, the ratio would be:

"\frac{P*R^2}{3P*R^2+P*R^3} = \frac{11}{37}"

By solving this equation we get rate of interest:

"R=\frac{4}{11}\approx36.36%" %